OpenStudy (anonymous):

Divide. Assume that all expressions are defined. x^2−25/2x^2+ 5x−12÷x^2−3x−10/x^2+ 9x + 20

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

\[\frac{ x^2-25 }{ 2x^2+5x-12}\div \frac{ x^2-3x-10 }{ x^2+9x+20 }\]

OpenStudy (anonymous):

First I would do the reciprocal

OpenStudy (anonymous):

which is

OpenStudy (anonymous):

the second part of the division, you flip it so you can multiply

OpenStudy (anonymous):

do you get me?

OpenStudy (anonymous):

yes i do, but i have never learned this and the lesson I am doing is not explaining this in full detail.

OpenStudy (anonymous):

ok so now you have to factor (x^2-25)

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (anonymous):

yes it is (x-5)(x+5)

OpenStudy (anonymous):

ok so keep that in the numerator and then factor the denominator 2x^2+5x-12

OpenStudy (anonymous):

i am confused by the 2x^2

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):



OpenStudy (anonymous):

Im sorry but i cant really explain how I got it?

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

Can you see it how i got it or no?

OpenStudy (anonymous):

a little

OpenStudy (anonymous):

I got the 3 and the 4 because you have to look for numbers that factor out to 12, like 3 * 4 and 6*2. and you have to start off with the 2x because you can't cut it in half to make 2x^2.

OpenStudy (anonymous):

do you get what im saying or no? you have to try out the factors to see which one works.

OpenStudy (anonymous):

okay i get it now thank you for explaining that

OpenStudy (anonymous):

okay so now we work on the other side. Factor x^2+9x+20. If you dont know how, I can tell you in detail how to factor this one?

OpenStudy (anonymous):

(x+5)(x+4)

OpenStudy (anonymous):

yass, ok so now factor x^2-3x-10.

OpenStudy (anonymous):

(x-5)(x+2)

OpenStudy (anonymous):

ok can you show what you all together?

OpenStudy (anonymous):

*what you have all together?

OpenStudy (anonymous):

\[\frac{ (x+5)(x-5) }{ (2x-3)(x+4) }\div\frac{ (x-5)(x+2) }{ (x+5)(x+4) }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

When dividing you have to flip the second division part, so you can multiply

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

I got is correct, now hou would I simplify this: \[\frac{-2x^2+2x }{ x^2-3x+2 }\]

OpenStudy (anonymous):

i get the bottom but not the top